# Quantum Optimization for the Maximum Cut Problem on a Superconducting Quantum Computer

### v1

@@ Acknowledgements.
where $\hat{X}_{i}$ is the Pauli operator on qubit $i$, $\hat{H}$ is the one-qubit Hadamard gate, the operator $\hat{\tilde{C}}$ is defined from the objective function, and $\{\gamma,\beta\}$ are real-valued angles. It is possible to obtain an analytic expression for the expectation value of the two-point correlation $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ between qubits $i$ and $j$, as a function of the parameters $\gamma$ and $\beta$. The expression was derived in App. B of Ref. [27] and we simply state the result:
@@
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}=$ $-\sin\bigl{(}2\beta\bigr{)}\cos\bigl{(}2\beta\bigr{)}\sin\bigl{(}\gamma\mathsf{W}_{ij}\bigr{)}\times\Biggl{[}\prod\nolimits_{k\neq i,j}\cos\bigl{(}\gamma\mathsf{W}_{ik}\bigr{)}+\prod\nolimits_{k\neq i,j}\cos\bigl{(}\gamma\mathsf{W}_{jk}\bigr{)}\Biggr{]}$
@@
$-\frac{\sin^{2}\bigl{(}2\beta\bigr{)}}{2}\Biggl{[}\prod\nolimits_{k\neq i,j}\cos\gamma\bigl{(}\mathsf{W}_{ik}+\mathsf{W}_{jk}\bigr{)}-\prod\nolimits_{k\neq i,j}\cos\gamma\bigl{(}\mathsf{W}_{jk}-\mathsf{W}_{ik}\bigr{)}\Biggr{]}.$ (S30)
@@
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}_{i\neq j}=-2\sin(2\beta)\cos(2\beta)\sin(\gamma)\cos^{2}(\gamma)\Biggl{\{}\mathsf{W}_{ij}+\frac{\tan\bigl{(}2\beta\bigr{)}\cos^{2}\bigl{(}\gamma\bigr{)}\sin\bigl{(}\gamma\bigr{)}}{2}\Bigl{[}\mathsf{N}_{ij}+\delta_{\mathsf{N}_{ij}3}\tan^{4}\bigl{(}\gamma\bigr{)}+\mathsf{W}_{ij}\mathsf{N}_{ij}\tan^{2}\bigl{(}\gamma\bigr{)}\Bigr{]}\Biggr{\}},$ (S31)
where $\mathsf{N}_{ij}=\sum_{k}\mathsf{W}_{ik}\mathsf{W}_{kj}$ and $\delta_{\mathsf{N}_{ij}3}$ is the Kronecker delta. The quantity $0\leq\mathsf{N}_{ij}\leq 3$ counts the number of times nodes $i$ and $j$ share a common nearest-neighbor $k$, independently of whether $i$ and $j$ are nearest-neighbors themselves. In Fig. S14 we show the local structures around nodes $i$ and $j$ of $3$-regular graphs and the corresponding values of $\mathsf{N}_{ij}$ and $\mathsf{W}_{ij}$.
Figure S14: Local structures appearing in $3$-regular graphs with respect to two nodes $i$ and $j$. (a) $\mathsf{N}_{ij}=\mathsf{W}_{ij}=0$ (b) $\mathsf{N}_{ij}=0$ and $\mathsf{W}_{ij}=1$ (c) $\mathsf{N}_{ij}=1$ and $\mathsf{W}_{ij}=0$ (d) $\mathsf{N}_{ij}=1$ and $\mathsf{W}_{ij}=1$ (e) $\mathsf{N}_{ij}=2$ and $\mathsf{W}_{ij}=0$ (f) $\mathsf{N}_{ij}=2$ and $\mathsf{W}_{ij}=1$ (g) $\mathsf{N}_{ij}=3$ and $\mathsf{W}_{ij}=0$.
S5.5 Commutator of Adjacency and Correlation Matrices
S5.5.1 Analytical Derivation Based on the QAOA at $p=1$
For $i\neq j$, the correlation matrix entries are simply the signed correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{i\neq j}$ in Eq. (S31) and $\mathsf{Z}_{i=j}=0$ otherwise. In the following, we show that for unit-weight $3$-regular graphs, the correlation and adjacency matrices commute in the large $N$ limit. As a result, the two matrices share the same eigenvectors, assuming that either matrix has distinct eigenvalues. As such, a relax-and-round approach on either of the two matrices will provide the same solution. The matrix resulting from the commutator $[\mathsf{W},\mathsf{Z}]$ has entries
$\bigl{[}\mathsf{W}\mathsf{Z}\bigr{]}_{ij}-\bigl{[}\mathsf{Z}\mathsf{W}\bigr{]}_{ij}=\sum\nolimits_{k}\mathsf{W}_{ik}\mathsf{Z}_{kj}-\sum\nolimits_{k}\mathsf{Z}_{ik}\mathsf{W}_{kj}.$ (S32)
There are four terms in $\mathsf{Z}_{ij}$ (see Eq. (S31)) which we consider in order:
@@
    The third term is $\propto\delta_{\mathsf{N}_{ij}3}$. Up to a prefactor depending on $\gamma$ and $\beta$, the commutator entries read $[\mathsf{W},\delta_{\mathsf{N}=3}]_{ij}=\sum_{k}\mathsf{W}_{ik}\delta_{\mathsf{N}_{kj}=3}-\sum_{k}\delta_{\mathsf{N}_{ik}=3}\mathsf{W}_{kj}$. According to Fig. S14g, if $\delta_{\mathsf{N}_{kj}=3}=1$ for two nodes $k$ and $j$, then, necessarily $\mathsf{W}_{ik}=1$ if and only if $\mathsf{W}_{jk}=\mathsf{W}_{kj}=0$ and $\mathsf{W}_{ij}=\mathsf{W}_{ji}=1$, and zero otherwise. As such, $[\mathsf{W},\delta_{\mathsf{N}=3}]_{ij}=0$ for all angles $\gamma$ and $\beta$.
@@
    The last term is $\propto\mathsf{W}_{ij}\mathsf{N}_{ij}$. Unlike the three other cases, it is not straightforward to show whether or not the adjacency matrix commutes with this term for arbitrary unit-weight $3$-regular graphs and angles. According to Fig. S14, this term is nonzero for local structures involving a triangle, where $\mathsf{W}_{ij}=1$ and $\mathsf{N}_{ij}\geq 1$. Up to a constant which depends on angles $\beta$ and $\gamma$, the elements $(i,j)$ of the commutator read $\sum_{k}\mathsf{W}_{ik}\mathsf{W}_{kj}\mathsf{N}_{kj}-\sum_{k}\mathsf{W}_{ik}\mathsf{N}_{ik}\mathsf{W}_{kj}$ where $\mathsf{N}_{ij}=\sum_{k}\mathsf{W}_{ik}\mathsf{W}_{kj}$. Given that asymptotically with the problem size $N$, the average number of triangles in a random $3$-regular graph is $4/3$ independently of $N$ (see Sec. S3.1), this term’s contribution to the overall commutator will be vanishingly small in the limit of large $N$.
@@
S5.5.2 Data
Figure S15: Norm of the commutator between the adjacency $\mathsf{W}$ and correlation $\mathsf{Z}^{(p)}$ matrices, as defined in Eq. (S33), as a function of the problem size $N$. Each data point is averaged over $1,000$ problem instances. Error bars indicate one standard deviation.
@@
$\Bigl{\|}\bigl{[}\mathsf{W},\mathsf{Z}^{(p)}\bigr{]}\Bigr{\|}_{2}=\Bigl{\|}\mathsf{W}\mathsf{Z}^{(p)}-\mathsf{Z}^{(p)}\mathsf{W}\Bigr{\|}_{2}=\sigma_{\textrm{max}},$ (S33)
where $\sigma_{\textrm{max}}$ is the largest singular value of the matrix resulting from the commutator. It is shown in Fig. S15 as a function of the problem size $N$ for $p=1$ and $p=2$. At $p=1$, we find that the norm of the commutator slowly decreases with $N$, in line with theoretical expectations. Although the norm of the commutator is not strictly zero at finite $N$, we verify in Sec. S5.1 that the solutions returned by the classical and quantum $(p=1)$ relax-and-round methods are the same. At $p=2$, the norm increases with $N$, in stark contrast to the behavior for $p=1$. Indeed, for $p>1$, the correlation matrix is not expected to commute with the adjacency matrix.
S5.6 Evaluating the Optimal Solution for Computing the Approximation Ratio
In order to compute the approximation ratio $\alpha$ (Eq. (S2)) of a solution $\boldsymbol{z}$, one needs to know the problem’s optimal solution, $\boldsymbol{z}_{\textrm{opt}}$. Here, we employ the best-in-class Burer-Monteiro classical heuristic [15], implemented in the library MQLib [56], and introduced in Sec. S2.1.4.
@@
For $N\leq 512$, we limited the total runtime to one second and found that each problem instance was solved about $100$ times, with a peak in the distribution of solutions at a minimum objective value $\tilde{C}$. Thus, we are quite confident in the optimality of the solution used for estimating the approximation ratio.
@@
S5.7 Performance Gain from the Greedy-Enhanced Quantum-Relax-and-Round Solver
We study the performance gain of the greedy-enhanced quantum relax-and-round solver (Sec. S2.2.4) over its non-enhanced version (Sec. S2.2.2). We implement two versions, where sign-flips are attempted on variables that are: (i) randomly selected; and (ii) guided by the non-rounded eigenvector entries. We define the numbers of variables visited as $fN$ for an $N$-variable problem with $f\in\mathbb{R}$ a control parameter.
Figure S16: Average distance to the perfect approximation ratio, as a function of the number of greedy steps. Results are presented for both the random and guided sign-flip strategies with problem sizes: (a) $N=32$, (b) $N=64$, (c) $N=128$, (d) $N=256$, (e) $N=512$, (f) $N=1024$, (g) $N=2048$, and (h) $N=4096$. Each data point is averaged over $1,000$ randomly generated problem instances. Error bars indicate one standard deviation. Number of steps are in units of $N$ i.e., $f$.
We show in Fig. S16 the average distance to the perfect approximation ratio as a function of $f$ for both sign-flipping strategies and various problem sizes $N$. We find that they both converge towards the same, higher, average approximation ratio as compared to the nonenhanced standard version. We observe that the guided version converges faster, but the random version asymptotically yields the same average approximation ratio. The convergence occurs with $f\sim O(1)$, independently of $N$. This means the random version visits all variables $O(1)$ times on average, thus enabling it to visit variables that the guided version identifies as being favorable, albeit with somewhat slower convergence.
@@
S5.8 Comparison of Various Combinatorial Optimization Solvers
Figure S17: Approximation ratio averaged over $1,000$ randomly generated $3$-regular graph problem instances, computed using a range of solvers for each problem size $N$. Solvers are: “QAOAp” (QAOA with $p$ layers); “GRDY” (a classical greedy solver); “RR” (a classical relax-and-round solver); “SAk” (simulated annealing with $k$ sweeps); “PTk” (parallel tempering with $k$ sweeps); “BMt” (the Burer-Monteiro solver with maximum runtime $t$ in milliseconds; and “QRR*p (greedy-enhanced quantum relax-and-round-solver based on the QAOA with $p$ layers). For the quantum solvers, “sim” and “exp” indicate emulated and experimental results, respectively. The approximation ratio is presented as a percentage below each solver, with the standard error in parentheses.
In this section, we benchmark the performance of QRR* against a suite of methods. The results are shown in Fig. S17. The comparison includes the QAOA (Sec.  S2.2.1), the most competitive classical solvers (simulated annealing and the Burer-Monteiro algorithm; see Sections S2.1.2 and S2.1.4, respectively), and the following:
@@
The classical greedy solver (Sec. S2.1.6) serves as a comparison point against quantum iterative solvers bearing similarities [108, 24, 109]. Fig. S17 shows the greedy solver is not competitive against QRR*. However, it should be noted that it performs better than the QAOA at $p=2$, highlighting the need to embed the QAOA into hybrid quantum-classical workflows (such as the QRR algorithm) to get the most out of it.
@@
The classical relax-and-round solver was compared against the classical and quantum Goemans-Williamson methods in Sec. S5.1. In the main text we also discussed how the QRR algorithm at $p=1$ is equivalent to the classical relax-and-round solver. Fig. S17 shows that the relax-and-round solver yields an average approximation ratio of about $97\%$, much higher than both the classical greedy algorithm and the QAOA.
@@
S5.9 Measuring the Runtime of Various Combinatorial Optimization Solvers
S5.9.1 Greedy-Enhanced Quantum Relax-and-Round Solver
The greedy-enhanced quantum relax-and-round solver (QRR*) of Sec. S2.2.4 is the reference quantum solver used throughout this work. Its runtime can be decomposed into four blocks: (i) runtime on the quantum processor for collecting bit strings; (ii) building the correlation matrix from the bit strings; (iii) computing the leading eigenvectors of the correlation matrix and sign-rounding them; and (iv) performing a greedy local search on the solution.
Important Introductory Remarks
The runtime on the quantum processor assumes that the circuit is built using a worst-case-scenario swap network on a hardware-native linear chain topology of qubits, which requires $Np$ layers of two-qubit gates [33]. An efficient compilation technique, such as the one described in Sec. S4.1.1.1, is not used.
The runtime on the quantum processor does not rely on the graph isomorphism described in Sec. S3.3, which was just a convenience for the purpose of this work to obtain statistics on thousands of problem instances. The runtime assumes that one has to execute a circuit for each of the nonzero entries of the correlation matrix.
Figure S18: (a) Time in seconds for executing and collecting $n_{\textrm{ex}}$ bit strings from QAOA circuits involving $N$ qubits and $p$ layers, compiled to a linear chain topology using a full swap network [33]. The plot shows the time to execute on the AnkaaTM-2 Quantum Processing Unit accessed through Rigetti Quantum Cloud ServicesTM (QCS), the current circuit duration and the basis for quantum timing in this paper (solid line), and a target execution time based on the best superconducting gate times in the literature. (b) Residuals between the model of Eq. S34 and the experimental data in panel (a).
Runtime on the Quantum Processor
We can measure the time it takes to execute a QAOA circuit compiled to a linear chain topology using a full swap network [33] on Rigetti Quantum Cloud ServicesTM (QCS) by instrumenting calls to the cloud service API as exposed by the Pyquil open-source software package. Using the AnkaaTM-2 Quantum Processing Unit (QPU) with $84$ qubits, we compiled $300$ test circuits across a range of $N$ (number of qubits), $p$ (number of QAOA layers), and $n_{\textrm{ex}}$ (number of bit strings collected). The test circuits implemented the QAOA, compiled to a linear chain topology using a full swap network [33]. The results are shown in Fig. S18a.
The QCS runtime can be modeled as
$t_{\textrm{ex}}=c_{0}+c_{N}N+c_{n_{\textrm{ex}}}n_{\textrm{ex}}+c_{N.p}Np+c_{N.n_{\textrm{ex}}}Nn_{\textrm{ex}}+c_{N.p.n_{\textrm{ex}}}Npn_{\textrm{ex}},$ (S34)
where each $c_{i}$ is a prefactor for a specific computational expense that might be improved. When we fit this model to the measured data using least-squares fitting over a linear residual, we obtain the coefficients of Tab. S4. The residual of the fit is illustrated in Fig. S18b.

### v2

@@ E.4.1 Analytical Derivation Based on the QAOA at $p=1$
At $p=1$, it is possible to obtain an analytic expression for the expectation value of the two-point correlation $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ between qubits $i$ and $j$, as a function of the parameters $\gamma_{1}\equiv\gamma$ and $\beta_{1}\equiv\beta$. The expression was derived in App. B of Ref. [29] and we simply state the result:
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}=$ $-\sin\bigl{(}2\beta\bigr{)}\cos\bigl{(}2\beta\bigr{)}\sin\bigl{(}\gamma\mathsf{W}_{ij}\bigr{)}$
@@
×$\times\Biggl{[}\prod\nolimits_{k\neq i,j}\cos\bigl{(}\gamma\mathsf{W}_{ik}\bigr{)}+\prod\nolimits_{k\neq i,j}\cos\bigl{(}\gamma\mathsf{W}_{jk}\bigr{)}\Biggr{]}$
@@
$-\frac{\sin^{2}\bigl{(}2\beta\bigr{)}}{2}\Biggl{[}\prod\nolimits_{k\neq i,j}\cos\gamma\bigl{(}\mathsf{W}_{ik}+\mathsf{W}_{jk}\bigr{)}$
$-\prod\nolimits_{k\neq i,j}\cos\gamma\bigl{(}\mathsf{W}_{jk}-\mathsf{W}_{ik}\bigr{)}\Biggr{]}.$ (32)
@@
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}_{i\neq j}=$ $-2\sin(2\beta)\cos(2\beta)\sin(\gamma)\cos^{2}(\gamma)$
×$\times\Biggl{\{}\mathsf{W}_{ij}+\frac{\tan\bigl{(}2\beta\bigr{)}\cos^{2}\bigl{(}\gamma\bigr{)}\sin\bigl{(}\gamma\bigr{)}}{2}$
×$\times\Bigl{[}\mathsf{N}_{ij}+\delta_{\mathsf{N}_{ij}3}\tan^{4}\bigl{(}\gamma\bigr{)}+\mathsf{W}_{ij}\mathsf{N}_{ij}\tan^{2}\bigl{(}\gamma\bigr{)}\Bigr{]}\Biggr{\}},$ (33)
where $\mathsf{N}_{ij}=\sum_{k}\mathsf{W}_{ik}\mathsf{W}_{kj}$ and $\delta_{\mathsf{N}_{ij}3}$ is the Kronecker delta. The quantity $0\leq\mathsf{N}_{ij}\leq 3$ counts the number of times nodes $i$ and $j$ share a common nearest-neighbor $k$, independently of whether $i$ and $j$ are nearest-neighbors themselves. In Fig. 18 we show the local structures around nodes $i$ and $j$ of $3$-regular graphs and the corresponding values of $\mathsf{N}_{ij}$ and $\mathsf{W}_{ij}$.
Figure 18: Local structures appearing in $3$-regular graphs with respect to two nodes $i$ and $j$. (a) $\mathsf{N}_{ij}=\mathsf{W}_{ij}=0$ (b) $\mathsf{N}_{ij}=0$ and $\mathsf{W}_{ij}=1$ (c) $\mathsf{N}_{ij}=1$ and $\mathsf{W}_{ij}=0$ (d) $\mathsf{N}_{ij}=1$ and $\mathsf{W}_{ij}=1$ (e) $\mathsf{N}_{ij}=2$ and $\mathsf{W}_{ij}=0$ (f) $\mathsf{N}_{ij}=2$ and $\mathsf{W}_{ij}=1$ (g) $\mathsf{N}_{ij}=3$ and $\mathsf{W}_{ij}=0$.
For $i\neq j$, the correlation matrix entries are simply the signed correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{i\neq j}$ in Eq. (33) and $\mathsf{Z}_{i=j}=0$ otherwise. In the following, we show that for unit-weight $3$-regular graphs, the correlation and adjacency matrices commute in the large $N$ limit. As a result, the two matrices share the same eigenvectors, assuming that either matrix has distinct eigenvalues. As such, a relax-and-round approach on either of the two matrices will provide the same solution. The matrix resulting from the commutator $[\mathsf{W},\mathsf{Z}]$ has entries
$\bigl{[}\mathsf{W}\mathsf{Z}\bigr{]}_{ij}-\bigl{[}\mathsf{Z}\mathsf{W}\bigr{]}_{ij}=\sum\nolimits_{k}\mathsf{W}_{ik}\mathsf{Z}_{kj}-\sum\nolimits_{k}\mathsf{Z}_{ik}\mathsf{W}_{kj}.$ (34)
There are four terms in $\mathsf{Z}_{ij}$ (see Eq. (33)) which we consider in order:
@@
    The third term is $\propto\delta_{\mathsf{N}_{ij}3}$. Up to a prefactor depending on $\gamma$ and $\beta$, the commutator entries read $[\mathsf{W},\delta_{\mathsf{N}=3}]_{ij}=\sum_{k}\mathsf{W}_{ik}\delta_{\mathsf{N}_{kj}=3}-\sum_{k}\delta_{\mathsf{N}_{ik}=3}\mathsf{W}_{kj}$. According to Fig. 18g, if $\delta_{\mathsf{N}_{kj}=3}=1$ for two nodes $k$ and $j$, then, necessarily $\mathsf{W}_{ik}=1$ if and only if $\mathsf{W}_{jk}=\mathsf{W}_{kj}=0$ and $\mathsf{W}_{ij}=\mathsf{W}_{ji}=1$, and zero otherwise. As such, $[\mathsf{W},\delta_{\mathsf{N}=3}]_{ij}=0$ for all angles $\gamma$ and $\beta$.
@@
    The last term is $\propto\mathsf{W}_{ij}\mathsf{N}_{ij}$. Unlike the three other cases, it is not straightforward to show whether or not the adjacency matrix commutes with this term for arbitrary unit-weight $3$-regular graphs and angles. According to Fig. 18, this term is nonzero for local structures involving a triangle, where $\mathsf{W}_{ij}=1$ and $\mathsf{N}_{ij}\geq 1$. Up to a constant which depends on angles $\beta$ and $\gamma$, the elements $(i,j)$ of the commutator read $\sum_{k}\mathsf{W}_{ik}\mathsf{W}_{kj}\mathsf{N}_{kj}-\sum_{k}\mathsf{W}_{ik}\mathsf{N}_{ik}\mathsf{W}_{kj}$ where $\mathsf{N}_{ij}=\sum_{k}\mathsf{W}_{ik}\mathsf{W}_{kj}$. Given that asymptotically with the problem size $N$, the average number of triangles in a random $3$-regular graph is $4/3$ independently of $N$ (see App. C.1), this term’s contribution to the overall commutator will be vanishingly small in the limit of large $N$.
@@
E.4.2 Numerical Data
Figure 19: Norm of the commutator between the adjacency $\mathsf{W}$ and correlation $\mathsf{Z}^{(p)}$ matrices, as defined in Eq. (35), as a function of the problem size $N$. Each data point is averaged over $1,000$ problem instances. Error bars indicate one standard deviation.
@@ E.4.2 Numerical Data
$\Bigl{\|}\bigl{[}\mathsf{W},\mathsf{Z}^{(p)}\bigr{]}\Bigr{\|}_{2}=\Bigl{\|}\mathsf{W}\mathsf{Z}^{(p)}-\mathsf{Z}^{(p)}\mathsf{W}\Bigr{\|}_{2}=\sigma_{\textrm{max}},$ (35)
where $\sigma_{\textrm{max}}$ is the largest singular value of the matrix resulting from the commutator. It is shown in Fig. 19 as a function of the problem size $N$ for $p=1$ and $p=2$. At $p=1$, we find that the norm of the commutator slowly decreases with $N$, in line with theoretical expectations. Although the norm of the commutator is not strictly zero at finite $N$, we verify that the solutions returned by the classical and quantum $(p=1)$ relax-and-round methods are the same (see App. E.4). At $p=2$, the norm increases with $N$, in stark contrast to the behavior for $p=1$. Indeed, for $p>1$, the correlation matrix is not expected to commute with the adjacency matrix.
E.5 Evaluating the Optimal Solution for Computing the Approximation Ratio
In order to compute the approximation ratio $\alpha$ (Eq. (5)) of a solution $\boldsymbol{z}$, one needs to know the problem’s optimal solution, $\boldsymbol{z}_{\textrm{opt}}$. Here, we employ the best-in-class Burer-Monteiro classical heuristic [16], implemented in the library MQLib [54], and introduced in App. B.1.4.
@@ E.5 Evaluating the Optimal Solution for Computing the Approximation Ratio
For $N\leq 512$, we limited the total run-time to one second and found that each problem instance was solved about $100$ times, with a peak in the distribution of solutions at a minimum objective value $\tilde{C}$. Thus, we are quite confident in the optimality of the solution used for estimating the approximation ratio.
@@
E.6 Performance Gain from the Greedy-Enhanced Quantum-Relax-and-Round Solver
We study the performance gain of the greedy-enhanced quantum relax-and-round solver (App. B.2.4) over its non-enhanced version (App. B.2.2). We implement two versions, where sign-flips are attempted on variables that are: (i) randomly selected; and (ii) guided by the non-rounded eigenvector entries. We define the numbers of variables visited as $fN$ for an $N$-variable problem with $f\in\mathbb{R}$ a control parameter.
Figure 20: Average distance to the perfect approximation ratio, as a function of the number of greedy steps. Results are presented for both the random and guided sign-flip strategies with problem sizes: (a) $N=32$, (b) $N=64$, (c) $N=128$, (d) $N=256$, (e) $N=512$, (f) $N=1024$, (g) $N=2048$, and (h) $N=4096$. Each data point is averaged over $1,000$ randomly generated problem instances. Error bars indicate one standard deviation. Number of steps are in units of $N$ i.e., $f$.
We show in Fig. 20 the average distance to the perfect approximation ratio as a function of $f$ for both sign-flipping strategies and various problem sizes $N$. We find that they both converge toward the same, higher, average approximation ratio as compared to the nonenhanced standard version. We observe that the guided version converges faster, but the random version asymptotically yields the same average approximation ratio. The convergence occurs with $f\sim O(1)$, independently of $N$. This means the random version visits all variables $O(1)$ times on average, thus enabling it to visit variables that the guided version identifies as being favorable, albeit with somewhat slower convergence.
@@ E.6 Performance Gain from the Greedy-Enhanced Quantum-Relax-and-Round Solver
E.7 Comparison of Various Combinatorial Optimization Solvers
Figure 21: Approximation ratio averaged over $1,000$ randomly generated $3$-regular graph problem instances, computed using a range of solvers for each problem size $N$. Solvers are: “QAOAp” (QAOA with $p$ layers); “GRDY” (a classical greedy solver); “RR” (a classical relax-and-round solver); “SAk” (simulated annealing with $k$ sweeps); “PTk” (parallel tempering with $k$ sweeps); “BMt” (the Burer-Monteiro solver with maximum run-time $t$ in milliseconds; and “QRR*p (greedy-enhanced quantum relax-and-round-solver based on the QAOA with $p$ layers). For the quantum solvers, “sim” and “exp” indicate emulated and experimental results, respectively. The approximation ratio is presented as a percentage below each solver, with the standard error in parentheses.
In this section, we benchmark the performance of QRR* against a suite of methods. The results are shown in Fig. 21. The comparison includes the QAOA (App.  B.2.1), the most competitive classical solvers (simulated annealing and the Burer-Monteiro algorithm; see Appendices B.1.2 and B.1.4, respectively), and the following:
@@ E.7 Comparison of Various Combinatorial Optimization Solvers
The classical greedy solver (App. B.1.6) serves as a comparison point against quantum iterative solvers bearing similarities [106, 58, 107]. Fig. 21 shows the greedy solver is not competitive against QRR*. However, it should be noted that it performs better than the QAOA at $p=2$, highlighting the need to embed the QAOA into hybrid quantum-classical workflows (such as the QRR algorithm) to get the most out of it.
@@
The classical relax-and-round solver was compared against the classical and quantum Goemans-Williamson methods in App. E.1. In the main text we also discussed how the QRR algorithm at $p=1$ is equivalent to the classical relax-and-round solver. Fig. 21 shows that the relax-and-round solver yields an average approximation ratio of about $97\%$, much higher than both the classical greedy algorithm and the QAOA.
@@
E.8 Measuring the Run-time of Various Combinatorial Optimization Solvers
E.8.1 Run-time of the Greedy-Enhanced Quantum Relax-and-Round Solver
The greedy-enhanced quantum relax-and-round solver (QRR*) is the reference quantum solver used throughout this work. Its run-time can be decomposed into four blocks: (i) run-time on the quantum processor for collecting bit strings; (ii) building the correlation matrix from the bit strings; (iii) computing the leading eigenvectors of the correlation matrix and sign-rounding them; and (iv) performing a greedy local search on the solution.
The run-time on the quantum processor assumes that the circuit is built using a worst-case-scenario swap network on a hardware-native linear chain topology of qubits, which requires $Np$ layers of two-qubit gates [35]. An efficient compilation technique, such as the one described in App. D.1.1, is not used. Moreover, the run-time on the quantum processor does not rely on the graph isomorphism described in App. C.3, which was just a convenience for the purpose of this work to obtain statistics on thousands of problem instances. The run-time assumes that one has to execute a circuit for each of the nonzero entries of the correlation matrix.
Figure 22: (a) Run-time for computing the correlation matrix made of $n_{\textrm{entries}}$ nonzero entries from $n_{\textrm{ex}}$ bit strings. The run-time is linear per entry and bit string, with the prefactor approximately $3.859(3)\times 10^{-10}$ per least-squares fitting. (b) Run-time for the eigendecomposition of the $N\times N$ adjacency ($\mathsf{W}$) and correlation matrices ($\mathsf{Z}$) at $p=1$ and $p=2$ using the Lanczos algorithm. Each data point is averaged over $20$ matrices. Units in both panels: seconds.
Run-time on the Quantum Processor—
We can measure the time it takes to execute a QAOA circuit compiled to a linear chain topology using a full swap network [35] on Rigetti Quantum Cloud ServicesTM (QCS) by instrumenting calls to the cloud service API as exposed by the Pyquil open-source software package. Using the AnkaaTM-2 Quantum Processing Unit (QPU) with $84$ qubits, we compiled $300$ test circuits across a range of $N$ (number of qubits), $p$ (number of QAOA layers), and $n_{\textrm{ex}}$ (number of bit strings collected). The test circuits implemented the QAOA, compiled to a linear chain topology using a full swap network [35]. The results are shown in Fig. 17a.
The QCS run-time can be modeled as
$t_{\textrm{ex}}=$ $c_{0}+c_{N}N+c_{n_{\textrm{ex}}}n_{\textrm{ex}}+c_{N.p}Np$
$+c_{N.n_{\textrm{ex}}}Nn_{\textrm{ex}}+c_{N.p.n_{\textrm{ex}}}Npn_{\textrm{ex}},$ (36)
where each $c_{i}$ is a prefactor for a specific computational expense that might be improved. When we fit this model to the measured data using least-squares fitting over a linear residual, we obtain the coefficients of Tab. 4 such that $t_{\textrm{ex}}$ is in seconds. The residual of the fit is illustrated in Fig. 17b.